The length of a rectangle is 10 in. more than twice its width. If the perimeter of the rectangle is 32 in., find the width of the rectangle.
Thanks so much!

Answer by checkley71(8403)

L=2W+10
32=2(2W+10)+2W
32=4W+20+2W
32=6W+20
6W=32-20
6W=12
W=12/6
W=2 ANSWER FOR THE WIDTH.
L=2*2+10
L=4+10
L=14 ANSWER FOR THE LENGTH.
PROOF
2*2+2*14=32
4+28=32
32=32